OpenStudy (anonymous):

What is the value of the y variable in the solution to the following system of equations? 3x + 2y = 6 2x + 3y = −6 A: 6 B: −3 C: 3 D: −6 *i got -3 but i want to make sure im right.

OpenStudy (amistre64):

it might be right, how did you obtain it?

OpenStudy (amistre64):

i dont get a -3

OpenStudy (jiteshmeghwal9):

2(3x+2y)=2(6) 3(2x+3y)=3(-6)

OpenStudy (anonymous):

well i did it some weird way , like my own way so thats why i wasn't sure if it was right

OpenStudy (amistre64):

if we know the way you tried it, we can hone it better

OpenStudy (anonymous):

6x + 4y = 12? 6x + 9y = -18? @jiteshmeghwal9

OpenStudy (jiteshmeghwal9):

i have make one of the variable equal to gt the value of 'y' :)

OpenStudy (amistre64):

equal but opposite is another strategy that eliminates an extra step out of that

OpenStudy (jiteshmeghwal9):

i deleted my reply before u see

OpenStudy (anonymous):

okay wait can you explain it to me again?

OpenStudy (amistre64):

lol ;)

OpenStudy (anonymous):

im soo confused....

OpenStudy (amistre64):

it would be better to work off of the method you tried on your own .. imo

mathslover (mathslover):

cool down friend amistre is here .. be happy and enjoy the learning @Rkb<3NH

OpenStudy (anonymous):

yea but i got 6x = 2 at the end

OpenStudy (amistre64):

im not concerned with the endl id like to see the process you did to correct it

OpenStudy (anonymous):

ok wait

OpenStudy (anonymous):

3x + 2y = 6 -3x -3x 2y = 6 - 3x 2x + 3y = −6 2x +(6-3X) 3y = −6 12x - 6x + 3y = -6 6x + 3y = -6 / 3 /3 6x = 2??

OpenStudy (amistre64):

so your trying to do substitution

OpenStudy (anonymous):

yea i tried elimation but to me it dident work out

OpenStudy (amistre64):

you should start out by solving for an "x" to find the "y" in order to replace the "x"; we substitute it with its y equivalent does that make sense?

OpenStudy (amistre64):

3x + 2y = 6 -2y -2y ----------- 3x = 6-2y x = (6-2y)/3

OpenStudy (anonymous):

yes but why did you divide by 3?

OpenStudy (amistre64):

because 3x does not equal x

OpenStudy (amistre64):

3x/3 = x

OpenStudy (amistre64):

3x = 6-2y /3 /3 ------------ x = (6-2y)/3

OpenStudy (anonymous):

so i will be 2x + (6 -2y)/3 +3y = -6 12x + 4y/3 + 3y = -6 is that good so far?

OpenStudy (anonymous):

4x +4/3y +3y = -6 i went wrong here, i dont knw

OpenStudy (anonymous):

what to do

OpenStudy (amistre64):

almost; lets clarify this "substitution" given that: x = (6-2y)/3 and 2x + 3y = -6 REPLACE (substitute) the equation of "x" in place of "x" \[2x + 3y = -6\] \[2\frac{(6-2y)}{3} + 3y = -6\] ^^^ we replaced "x"

OpenStudy (anonymous):

12x - 4y/3 + 3y = -6 isn't that what i jut got because i cant divide 4/3

OpenStudy (amistre64):

there shouldnt be any "x" left to play with in this equation

OpenStudy (amistre64):

(12 - 4y)/3 + 3y = -6

OpenStudy (anonymous):

4y - 4/3y + 3y = -6?

OpenStudy (amistre64):

lets work this slowly (12 - 4y)/3 + 3y = -6\[\frac{12 - 4y}{3} + 3y = -6\]\[\frac{12}{3} - \frac{4y}{3} + 3y = -6\]\[4 - \frac{4}{3}y + 3y = -6\]-4 -4 -----------------\[- \frac{4}{3}y + 3y = -10\] good so far?

OpenStudy (jiteshmeghwal9):

6x+4y=12 6x=12+4y x=12+4y/6

mathslover (mathslover):

@jiteshmeghwal9 just a request to you : let amistre64 finish his explanation to the user .. please

OpenStudy (jiteshmeghwal9):

k!

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

and jite, your mathing; 6x+4y=12 6x 12+4y ; is off a little just saying :)

OpenStudy (anonymous):

but 4/3y + 3y is 4.33333333...

OpenStudy (amistre64):

Rkb, good. as long as its making sense to far, lets continue; without worrying about turning fractions into decimals .... \[- \frac{4}{3}y + 3y = -10\] \[\left(3-\frac{4}{3}\right)y = -10\] \[\left(\frac{9}{3}-\frac{4}{3}\right)y = -10\] \[\frac{5}{3}y = -10\] still good?

OpenStudy (anonymous):

ok i get it now so it will be 3 * 5/3 = -10 * 3 5 = 30? good so far?

OpenStudy (amistre64):

thats pretty good, dont forget the negative from that 10 tho 3 * 5/3 y = -10 * 3 5y = -30

OpenStudy (anonymous):

oh okay: y = -6!!??(:

OpenStudy (amistre64):

correct :) y = -6

OpenStudy (jiteshmeghwal9):

cool math @amistre64 :)

OpenStudy (amistre64):

thanx; elimination is quicker if you can work your head around it; but subbing is just as fine if your more comfortable with it

OpenStudy (jiteshmeghwal9):

yes! i was going through elimination method ut @mathslover saide to me that we should see ur explanation first :)

OpenStudy (anonymous):

thank you soo much @amistre64 (: And thank you tooo @jiteshmeghwal9 (: im glad you didn't just give me the answer you you helped me out(:

OpenStudy (amistre64):

3x + 2y = 6 2x + 3y = −6 3x + 2y = 6 ; *2 2x + 3y = −6 ; *-3 6x + 4y = 12 -6x -9y = 18 ; combine them -------------- -5y = 30 /-5 /-5 ----------- y = -6

OpenStudy (jiteshmeghwal9):

\[\Huge{\color{red}{you \space are \space welcome :) }}\]